# MgNO: Efficient Parameterization of Linear Operators via Multigrid

### v1

@@ Abstract
In this work, we propose a concise neural operator architecture for operator learning. Drawing an analogy with a conventional fully connected neural network, we define the neural operator as follows: the output of the $i$-th neuron in a nonlinear operator layer is defined by $\mathcal{O}_{i}(u)=\sigma\left(\sum_{j}\mathcal{W}_{ij}u+\mathcal{B}_{ij}\right)$. Here, $\mathcal{W}_{ij}$ denotes the bounded linear operator connecting $j$-th input neuron to $i$-th output neuron, and the bias $\mathcal{B}_{ij}$ takes the form of a function rather than a scalar. Given its new universal approximation property, the efficient parameterization of the bounded linear operators between two neurons (Banach spaces) plays a critical role. As a result, we introduce MgNO, utilizing multigrid structures to parameterize these linear operators between neurons. This approach offers both mathematical rigor and practical expressivity. Additionally, MgNO obviates the need for conventional lifting and projecting operators typically required in previous neural operators. Moreover, it seamlessly accommodates diverse boundary conditions. Our empirical observations reveal that MgNO exhibits superior ease of training compared to other CNN-based models, while also displaying a reduced susceptibility to overfitting when contrasted with spectral-type neural operators. We demonstrate the efficiency and accuracy of our method with consistently state-of-the-art performance on different types of partial differential equations (PDEs).
@@ 1 Introduction
Partial differential equation (PDE) models are ubiquitous in physics, engineering, and other disciplines. Many scientific and engineering fields rely on solving PDEs, such as optimizing airfoil shapes for better airflow, predicting weather patterns by simulating the atmosphere, and testing the strength of structures in civil engineering. Tremendous efforts have been made to solve various PDEs arising from different areas. Solving PDEs usually requires designing numerical methods that are tailored to the specific problem and depend on the insights of the model. However, deep learning models have shown great promise in solving PDEs in a more general and efficient way. Recently, several novel methods, including Fourier neural operator (FNO) [23], Galerkin transformer (GT) [2], deep operator network (DeepONet) [25], and convolutional neural operators [33], have been developed to directly learn the operator (mapping) between infinite-dimensional parameter and solution spaces of PDEs. These methods leverage advanced architectures, such as Fourier convolution and self-attentions, to handle a variety of input parameters and achieve high performance for forward and inverse PDE problems. The methodology has been applied to biomechanical engineering and weather forecasting [43, 31]. Although deep neural operators have shown great potential to learn complex patterns and relationships from data, they still have some limitations compared to the classical numerical methods, such as lower accuracy and less flexibility to complex domain and boundary conditions. To address these challenges, we investigate the intrinsic properties of PDE-governed tasks and design the neural operator architecture based on multigrid methods, one of the most commonly used and efficient classical numerical algorithms.
@@ 2 Background and related work
In recent times, substantial efforts have been dedicated to the development of neural networks or operators tailored for solving partial differential equations (PDEs). [25] introduced the DeepONet, employing a branch-trunk architecture grounded in the universal approximation theorem in [3]. FNO parameterizes the global convolutional operators using a fast Fourier transform. Furthermore, geo-FNO [22] addresses tasks involving intricate geometries, such as point clouds, by transforming data into a latent uniform mesh and back. U-NO [32] improves FNO with a U-shaped architecture. The latent spectral model (LSM) [39] tried to identify the latent space and represent the operator in latent space. In addition, MWT [9] introduces a multiwavelet-based operator by parametrizing the integral operators using a fast wavelet transform. Then, Cao explored the self-attention mechanism [38] and introduced a Galerkin-type attention mechanism with linear complexity for solving PDEs. More recently, a convolution-based architecture has been employed as a neural operator, as presented in [33]. Among the notable approaches, spectral-type neural operators have gained attention for their global operation characteristics. The non-local operators have shown very promising results. Mathematical viewpoints provide universal approximation property for such operators [18, 20]. However, their inherent limitation lies in their incapability to effectively handle boundary conditions. Methods like FNO, MWT, UNO, and LSM, rely on encoding boundary information within the input data, making them reliant on specific data representations. These methods tend to learn low-frequency component [24], which is not a surprise since high-frequency modes are often truncated for the spectral type method. In these methods, convolutions in the neural operator construction are parametrized by Fourier, or wavelet transforms. However, Fourier or wavelet-based methods are not always appropriate for solving PDEs because of boundary conditions and aliasing errors. In contrast, local and geometrical neural networks, such as ResNet, DilResNet, UNet [14, 15, 34], offer greater flexibility in managing diverse boundary conditions. Nevertheless, these approaches were not originally designed for solving PDEs and often necessitated a substantial number of parameters to achieve high accuracy. Furthermore, they tend to act as high-pass filters, focusing primarily on local features like edges, surfaces, and textures, which limits their ability to capture long-distance relationships crucial in various physical phenomena. Large kernel convolution could mitigate the issue, but the design of large kernels requires hand-crafted since it is often hard to train [6].
@@
Multigrid methods [10, 40, 37] stand out as some of the most efficient numerical approaches in scientific computing, especially when tackling elliptic PDEs. Intriguingly, within the deep learning community, multigrid methods first received mention in the original ResNet paper [14], where the authors cited the methods as a core rationale behind the utility of residuals. Subsequently, the authors in [12, 13] established deeper and more extensive links between multigrid methods and ResNet, giving rise to what is now known as MgNet. The multigrid methodology, along with MgNet’s principles, have recently found applications in solving numerical PDEs [4] and time series forecasting problems [44]. However, it is not clear how to apply multigrid to operator learning appropriately in these previous works.
@@ 3.2 Deep neural operators from $\mathcal{X}$ to $\mathcal{Y}$
We note that a standard hidden layer in GNO [19] or FNO [23] is given by
@@
From equation 3, a notable feature of our framework is the elimination of the commonly used lifting layer in the first layer and the projection layer in the last layer, which are prevalent in most prior neural operators [19, 22]. It’s important to highlight that these two artificial layers are considered crucial for demonstrating the approximation capabilities across a wide range of neural operators, as evidenced in [20].
@@ 4 Parametrization of $\mathcal{W}^{\ell}_{ij}$ and MgNO
For instance, when $d=2$, the input is discretized as $\bm{u}(x)=\sum_{i,j=1}^{d}u_{ij}\phi_{i,j}(x)\mapsto u\in\mathbb{R}^{d\times d}$, where $d=\frac{1}{h}$. Consequently, for any $\mathcal{W}^{\ell}_{ij}\in\mathcal{L}\left(\mathcal{Y},\mathcal{Y}\right)$, the dimensionality of the space $\mathcal{L}\left(\mathcal{Y},\mathcal{Y}\right)$ becomes $d^{2}\times d^{2}$. This dimensionality is challenging to parameterize directly for large values of $d$. Most existing neural operators, such as those in [19, 20], primarily offer low-rank approximations of the kernel function in the spectral or wavelet domain. In this work, we advocate for using multigrid structures to directly parameterize within the spatial domain.
@@ 4.1 Multigrid methods for discrete elliptic PDEs with boundary conditions
First, we offer a concise and practical overview of the rationale and methodology behind using multigrid techniques to parameterize Green’s kernel functions. Consider the elliptic PDEs given by $\mathcal{L}u(x)=f(x)$ defined over the domain $\Omega=(0,1)^{2}$ and subject to Dirichlet, Neumann, or periodic boundary conditions. Employing a linear FEM discretization with a mesh size defined as $h=\frac{1}{d}$, the discretized system can be expressed as:
@@
where $u,f\in\mathbb{R}^{d\times d}$. Here, $\ast$ represents the standard convolutional operation for a single channel, complemented by specific padding schemes determined by the boundary conditions. The kernel $A$, of dimensions $3\times 3$, is dictated by the elliptic operator $\mathcal{L}$ in conjunction with the linear FEM. Consequently, the inverse operation of $A\ast$ corresponds to the discrete Green’s kernel function associated with $\mathcal{L}$ under a linear FEM framework. As demonstrated in [12, 13], the V-cycle multigrid approach for solving equation 6 can be precisely represented as a conventional convolutional neural network with one-channel.
@@
    Subsequently, we apply the smoothing iteration as in  equation 7 to derive the correction from the coarser level. The correction is then prolonged from the coarser to the finer level using a de-convolution operation with a stride of 2 (acting as the transpose of the restriction operation). After this prolongation, one can either opt to prolong further to an even coarser level (known as the Backslash-cycle) or proceed with post-smoothing followed by prolongation (referred to as the V-cycle).
Let’s represent the linear operator defined by the aforementioned V-cycle multigrid operator as $\mathcal{W}_{Mg}$. The convergence result presented in [41]
@@
demonstrates the uniform approximation capabilities of $\mathcal{W}_{Mg}$ relative to the inverse of $A\ast$, which corresponds to the Green’s kernel function associated with the elliptic operator $\mathcal{L}$. Here, $c$ is a constant that is independent of the mesh size $h$, and $\|u\|^{2}_{A}=(u,A\ast u)_{L^{2}(\Omega)}$ denotes the energy norm.
From $\mathcal{W}^{\ell}_{ij}$ to $\mathcal{W}^{\ell}$: Balancing Expressivity and Complexity
We observe that directly using a one-channel $\mathcal{W}_{Mg}$ to parameterize each $\mathcal{W}^{\ell}_{ij}$ in equation 3 individually may lead to expressivity limitations. Efficiently parameterizing $\mathcal{W}^{\ell}$ for multi-neuron (or multi-channel) inputs remains a challenging task. In this study, drawing inspiration from the multi-channel MgNet presented in [12, 44], we advocate for a multi-channel multigrid approach to handle linear operators with multi-channel inputs and outputs. During implementation, it suffices to design all $A^{\ell}$, $B^{\ell,i}$, restriction, and prolongation kernels as multi-channel convolutional kernels. A detailed exposition of the multi-channel multigrid in the convolutional context is available in Appendix B, as described in Algorithm 23.
Two primary motivations drive our choice of this strategy. Firstly, as established in [11], channels in deep CNNs, for 2D inputs, function analogously to neurons in terms of their universal approximation capabilities. Consequently, augmenting the number of channels in multigrid enhances the expressiveness of $\mathcal{W}_{Mg}$. From a practical standpoint, the multi-channel CNN variant of $\mathcal{W}_{Mg}$ possesses a parameter count on the order of $\mathcal{O}\left(\log(d)n^{2}\right)$.
Figure 1: Overview of $\mathcal{W}_{Mg}$ using a multi-channel V-cycle multigrid framework. In equation 3, by substituting $\mathcal{W}^{\ell}\bm{h}^{\ell-1}$ with $\mathcal{W}_{Mg}^{\ell}\bm{h}^{\ell-1}$, we derive our MgNO as presented in equation 10. The term $\mathcal{W}_{Mg}^{\ell}\bm{h}^{\ell-1}$ represents the multi-channel V-cycle multigrid in Algorithm 23, using the multi-channel input $\bm{h}^{\ell-1}$. It’s important to note that Algorithm 23 does not include any nonlinear activation.
@@ 4.2 Architecture of MgNO
$\begin{cases}&\bm{h}^{0}=\bm{u}\in\mathcal{X},\\ &\bm{h}^{\ell}(\bm{u})=\bm{\sigma}\left(\mathcal{W}^{\ell}_{Mg}\bm{h}^{\ell-1}(\bm{u})+\bm{B}^{\ell}\bm{h}^{\ell-1}(\bm{u})+b^{\ell}\bm{1}\right)\in\mathcal{Y}^{n},\quad\ell=1:L,\\ &\bm{v}=\widetilde{\mathcal{G}}_{\theta}(\bm{u})=\mathcal{W}^{L+1}(\bm{h}^{L}(\bm{u}))\in\mathcal{Y}.\end{cases}$ (10)
Here, $\mathcal{W}^{\ell}_{Mg}\in\mathcal{L}\left(\mathcal{Y}^{n},\mathcal{Y}^{n}\right)$ represents the multi-channel multigrid as previously defined and detailed in Algorithm 23. The function $\sigma$ is the point-wise GELU activation, with $\bm{B}^{\ell}\in\mathbb{R}^{n\times n}$ and $b^{\ell}\in\mathbb{R}^{n}$. Notably, the total number of parameters in $\widetilde{\mathcal{G}}_{\theta}(\bm{u})$ is $\mathcal{O}(\log(d)n^{2}L)$, and its inference complexity stands at $\mathcal{O}(\log(d)d^{2}n^{2}L)$.
@@
It is imperative to highlight that MgNO is adept at accommodating the boundary conditions of various PDEs. Leveraging the inherent relationship between convolutions and multigrid, the output of $\mathcal{W}_{Mg}^{\ell}$ can seamlessly satisfy Dirichlet, Neumann, or periodic boundary conditions. Consequently, $\widetilde{\mathcal{G}}_{\theta}(u)$ can maintain these boundary conditions without any training phase, provided by taking $\bm{B}^{1}=0$ in equation 10 and ensuring $\bm{1}(x)$ satisfies the same boundary conditions. One can refer to Table 4 in E for specific convolution configuration for specific boundary conditions.

### v2

@@ Abstract
In this work, we propose a concise neural operator architecture for operator learning. Drawing an analogy with a conventional fully connected neural network, we define the neural operator as follows: the output of the $i$-th neuron in a nonlinear operator layer is defined by $\mathcal{O}_{i}(u)=\sigma\left(\sum_{j}\mathcal{W}_{ij}u+\mathcal{B}_{ij}\right)$. Here, $\mathcal{W}_{ij}$ denotes the bounded linear operator connecting $j$-th input neuron to $i$-th output neuron, and the bias $\mathcal{B}_{ij}$ takes the form of a function rather than a scalar. Given its new universal approximation property, the efficient parameterization of the bounded linear operators between two neurons (Banach spaces) plays a critical role. As a result, we introduce MgNO, utilizing multigrid structures to parameterize these linear operators between neurons. This approach offers both mathematical rigor and practical expressivity. Additionally, MgNO obviates the need for conventional lifting and projecting operators typically required in previous neural operators. Moreover, it seamlessly accommodates diverse boundary conditions. Our empirical observations reveal that MgNO exhibits superior ease of training compared to CNN-based models, while also displaying a reduced susceptibility to overfitting when contrasted with spectral-type neural operators. We demonstrate the efficiency and accuracy of our method with consistently state-of-the-art performance on different types of partial differential equations (PDEs).
@@ 1 Introduction
Partial differential equation (PDE) models are ubiquitous in physics, engineering, and other disciplines. Many scientific and engineering fields rely on solving PDEs, such as optimizing airfoil shapes for better airflow, predicting weather patterns by simulating the atmosphere, and testing the strength of structures in civil engineering. Tremendous efforts have been made to solve various PDEs arising from different areas. Solving PDEs usually requires designing numerical methods that are tailored to the specific problem and depend on the insights of the model. However, deep learning models have shown great promise in solving PDEs in a more general and efficient way. Recently, several novel methods, including Fourier neural operator (FNO) (Li et al., 2020), Galerkin transformer (GT) (Cao, 2021), deep operator network (DeepONet) (Lu et al., 2021), and convolutional neural operators Raonic et al. (2023), have been developed to directly learn the operator (mapping) between infinite-dimensional parameter and solution spaces of PDEs. These methods leverage advanced architectures, such as Fourier convolution and self-attentions, to handle a variety of input parameters and achieve high performance for forward and inverse PDE problems. The methodology has been applied to biomechanical engineering and weather forecasting (You et al., 2022; Pathak et al., 2022). Although deep neural operators have shown great potential to learn complex patterns and relationships from data, they still have some limitations compared to the classical numerical methods, such as lower accuracy and less flexibility to complex domain and boundary conditions. To address these challenges, we investigate the intrinsic properties of PDE-governed tasks and design the neural operator architecture based on multigrid methods, one of the most commonly used and efficient classical numerical algorithms.
@@ 2 Background and related work
In recent times, substantial efforts have been dedicated to the development of neural networks or operators tailored for solving partial differential equations (PDEs). Lu et al. (2021) introduced the DeepONet, employing a branch-trunk architecture grounded in the universal approximation theorem in Chen et al. (1996). FNO parameterizes the global convolutional operators using a fast Fourier transform. Furthermore, geo-FNO Li et al. (2022) addresses tasks involving intricate geometries, such as point clouds, by transforming data into a latent uniform mesh and back. U-NO Rahman et al. (2022) improves FNO with a U-shaped architecture. The latent spectral model (LSM) (Wu et al., 2023) tried to identify the latent space and represent the operator in latent space. In addition, MWT Gupta et al. (2021) introduces a multiwavelet-based operator by parametrizing the integral operators using a fast wavelet transform. Then, Cao explored the self-attention mechanism Vaswani et al. (2017) and introduced a Galerkin-type attention mechanism with linear complexity for solving PDEs. More recently, a convolution-based architecture has been employed as a neural operator, as presented in Raonic et al. (2023). Among the notable approaches, spectral-type neural operators have gained attention for their global operation characteristics. The non-local operators have shown very promising results. Mathematical viewpoints provide universal approximation property for such operators (Kovachki et al., 2021; Lanthaler et al., 2023). However, their inherent limitation lies in their incapability to effectively handle boundary conditions. Methods like FNO, MWT, UNO, and LSM, rely on encoding boundary information within the input data, making them reliant on specific data representations. These methods tend to learn low-frequency component (Liu et al., 2022), which is not a surprise since high-frequency modes are often truncated for the spectral type method. In these methods, convolutions in the neural operator construction are parametrized by Fourier, or wavelet transforms. However, Fourier or wavelet-based methods are not always appropriate for solving PDEs because of boundary conditions and aliasing errors. In contrast, local and geometrical neural networks, such as ResNet, DilResNet, UNet He et al. (2016a; b); Ronneberger et al. (2015), offer greater flexibility in managing diverse boundary conditions. Nevertheless, these approaches were not originally designed for solving PDEs and often necessitated a substantial number of parameters to achieve high accuracy. Furthermore, they tend to act as high-pass filters, focusing primarily on local features like edges, surfaces, and textures, which limits their ability to capture long-distance relationships crucial in various physical phenomena. Large kernel convolution could mitigate the issue, but the design of large kernels requires hand-crafted since it is often hard to train (Ding et al., 2022).
@@
Multigrid methods Hackbusch (2013); Xu (1989); Trottenberg et al. (2000) stand out as some of the most efficient numerical approaches in scientific computing, especially when tackling elliptic PDEs. Intriguingly, within the deep learning community, multigrid methods first received mention in the original ResNet paper He et al. (2016a), where the authors cited the methods as a core rationale behind the utility of residuals. Subsequently, the authors in He & Xu (2019); He et al. (2023) established deeper and more extensive links between multigrid methods and ResNet, giving rise to what is now known as MgNet. The research rigorously demonstrated that the linear V-cycle multigrid structure for the Poisson equation, can be represented as a convolutional neural network, despite their experimental focus on nonlinear multigrid with single-cycle structure for vision-related tasks. The original MgNet and its subsequent adaptations in the context of numerical PDEs Chen et al. (2022) and forecasting scenarios Zhu et al. (2023), have not provided a simple but effective solution in the realm of operator learning. In this work, we aim to fundamentally integrate multigrid methodologies with operator learning based on a very general framework.
@@ 3.2 Deep neural operators from $\mathcal{X}$ to $\mathcal{Y}$
We note that a standard hidden layer in GNO Kovachki et al. (2023) or FNO Li et al. (2020) is given by
@@
From equation 3, a notable feature of our framework is the elimination of the commonly used lifting layer in the first layer and the projection layer in the last layer, which are prevalent in most prior neural operators Kovachki et al. (2023); Li et al. (2022). It’s important to highlight that these two artificial layers are considered crucial for demonstrating the approximation capabilities across a wide range of neural operators, as evidenced in Lanthaler et al. (2023).
@@ 4 Parametrization of $\mathcal{W}^{\ell}_{ij}$ and MgNO
For instance, when $d=2$, the input is discretized as $\bm{u}(x)=\sum_{i,j=1}^{d}u_{ij}\phi_{i,j}(x)\mapsto u\in\mathbb{R}^{d\times d}$, where $d=\frac{1}{h}$. Consequently, for any $\mathcal{W}^{\ell}_{ij}\in\mathcal{L}\left(\mathcal{Y},\mathcal{Y}\right)$, the dimensionality of the space $\mathcal{L}\left(\mathcal{Y},\mathcal{Y}\right)$ becomes $d^{2}\times d^{2}$. This dimensionality is challenging to parameterize directly for large values of $d$. Most existing neural operators, such as those in Kovachki et al. (2023); Lanthaler et al. (2023), primarily offer low-rank approximations of the kernel function in the spectral or wavelet domain. In this work, we advocate for using multigrid structures to directly parameterize within the spatial domain.
@@ 4.1 Multigrid methods for discrete elliptic PDEs with boundary conditions
First, we offer a concise and practical overview of the rationale and methodology behind using multigrid techniques to parameterize Green’s functions. Consider the elliptic PDEs given by $\mathcal{L}u(x)=f(x)$ defined over the domain $\Omega=(0,1)^{2}$ and subject to Dirichlet, Neumann, or periodic boundary conditions. Employing a linear FEM discretization with a mesh size defined as $h=\frac{1}{d}$, the discretized system can be expressed as:
@@
where $u,f\in\mathbb{R}^{d\times d}$. Here, $\ast$ represents the standard convolutional operation for a single channel, complemented by specific padding schemes determined by the boundary conditions. The kernel $A$, of dimensions $3\times 3$, is dictated by the elliptic operator $\mathcal{L}$ in conjunction with the linear FEM. Consequently, the inverse operation of $A\ast$ corresponds to the discrete Green’s function associated with $\mathcal{L}$ under a linear FEM framework. As demonstrated in  He & Xu (2019); He et al. (2021), the V-cycle multigrid approach for solving equation 6 can be precisely represented as a conventional convolutional neural network with one-channel.
@@
    Subsequently, we apply the smoothing iteration as in  equation 7 to derive the correction from the coarser level. The correction is then prolonged from the coarser to the finer level using a de-convolution operation $P_{\ell+1}^{\ell}$ with a stride of 2 (acting as the transpose of the restriction operation). After this prolongation, one can either opt to prolong further to an even coarser level (known as the Backslash-cycle) or proceed with post-smoothing followed by prolongation (referred to as the V-cycle).
Let’s represent the linear operator defined by the aforementioned V-cycle multigrid operator as $\mathcal{W}_{Mg}$. The convergence result presented in Xu & Zikatanov (2002)
@@
demonstrates the uniform approximation capabilities of $\mathcal{W}_{Mg}$ relative to the inverse of $A\ast$, which corresponds to the Green’s function associated with the elliptic operator $\mathcal{L}$. Here, $c$ is a constant that is independent of the mesh size $h$, and $\|u\|^{2}_{A}=(u,A\ast u)_{L^{2}(\Omega)}$ denotes the energy norm. Please refer the Section C in Appendix for more details regarding the approximation property. We quantify the constant $c$ in equation 9 numerically using a concrete example.
Figure 1: Overview of $\mathcal{W}_{Mg}$ using a multi-channel V-cycle multigrid framework.
@@ 4.2 Architecture of MgNO
$\begin{cases}&\bm{h}^{0}=\bm{u}\in\mathcal{X},\\ &\bm{h}^{\ell}(\bm{u})=\bm{\sigma}\left(\mathcal{W}^{\ell}_{Mg}\bm{h}^{\ell-1}(\bm{u})+\bm{B}^{\ell}\bm{h}^{\ell-1}(\bm{u})+b^{\ell}\bm{1}\right)\in\mathcal{Y}^{n},\quad\ell=1:L,\\ &\bm{v}=\widetilde{\mathcal{G}}_{\theta}(\bm{u})=\mathcal{W}^{L+1}_{Mg}(\bm{h}^{L}(\bm{u}))\in\mathcal{Y}.\end{cases}$ (10)
Here, $\mathcal{W}^{1}_{Mg}\in\mathcal{L}\left(\mathcal{X},\mathcal{Y}^{n}\right)$, $\mathcal{W}^{\ell}_{Mg}\in\mathcal{L}\left(\mathcal{Y}^{n},\mathcal{Y}^{n}\right)$, for $\ell=2:J$ and $W^{L+1}_{Mg}\in\mathcal{L}\left(\mathcal{Y}^{n},\mathcal{Y}\right)$ are multi-channel linear operators that process multi-channel input $\bm{h}^{\ell-1}$ and output $\mathcal{W}^{\ell}_{Mg}(\bm{h}^{\ell-1})$. The function $\sigma$ is the point-wise GELU activation, with $\bm{B}^{\ell}\in\mathbb{R}^{n\times n}$ and $b^{\ell}\in\mathbb{R}^{n}$.
We observe that directly using a one-channel $\mathcal{W}_{Mg}$ to parameterize each $\mathcal{W}^{\ell}_{ij}$ in equation 3 individually lead to expressivity and efficiency limitations. It is advantageous to design all $A^{\ell}$, $B^{\ell,i}$, restriction operators $R_{\ell}^{\ell+1}$, and prolongation operators $P_{\ell+1}^{\ell}$, for $\ell=1:J$, as multi-channel convolutional kernels. A detailed exposition of the multi-channel multigrid in the convolutional context is available in Algorithm 23 Appendix B and figure 1. Two primary motivations drive our choice of this strategy. Firstly, as established in He et al. (2022), channels in deep CNNs, for 2D inputs, function analogously to neurons in terms of their universal approximation capabilities. Consequently, augmenting the number of channels in multigrid enhances the expressiveness of $\mathcal{W}_{Mg}$. Secondly, this approach aligns with classical multigrid methods applied to solve elliptic partial differential equation system, including linear elasticity problems in 2D or 3D scenarios. In such cases, the V-cycle multigrid methods can be naturally represented as linear MgNO with multi-channel inputs. From a practical standpoint, the multi-channel variant of $\mathcal{W}_{Mg}$ possesses a parameter count on the order of $\mathcal{O}\left(\log(d)n^{2}\right)$ and complexity on the order of $\mathcal{O}(d^{2}n^{2})$. Ultimately, by substituting $\mathcal{W}^{\ell}$ in equation 3, with $\mathcal{W}_{Mg}^{\ell}$, we derive our MgNO as presented in equation 10. It’s important to note that $\mathcal{W}^{\ell}_{Mg}$ does not include any nonlinear activation.
@@
It is imperative to highlight that MgNO is adept at accommodating the boundary conditions of various PDEs. Leveraging the inherent relationship between convolutions and multigrid, the output of $\mathcal{W}_{Mg}^{\ell}$ can seamlessly satisfy Dirichlet, Neumann, or periodic boundary conditions. Consequently, $\widetilde{\mathcal{G}}_{\theta}(u)$ can maintain these boundary conditions without any training phase, provided by taking $\bm{B}^{1}=0$ in equation 10 and ensuring $\bm{1}(x)$ satisfies the same boundary conditions. One can refer to Table 4 in Section F for specific convolution configuration for specific boundary conditions, and also Section C in Appendix for an concrete example of integrating boundary conditions into neural operators.